Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: In A configuration (ending 12 Sep); Correlator hardware (64 BlB) available. Realized improvements in peak data rates. New OPT release today. Maintenance day today (due to fixed-date daytime observing tomorrow).
  • Key dates:
    • 08 Jun - 12 Sep: A configuration (move from 31 May - 08 Jun).
  • Staff Travel:
    • RP: 09 Aug - 21 Aug (URSI)
    • FO: 12 Aug - 12 Sep
    • RP: 10 Sep - 16 Sep (Brazil)
  • Visitors:
    • Michiel Brentjens: 11 Jun - 11 Nov (Holography; characterizing polarized beams)
    • Dave Roberts: 10 Aug - 20 Aug (SS433 Demo science)

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • Antennas: EA25 out; EA23 on MP.
    • Receivers: EA23 gets outfitted last week in August.
      • Perley: RF Switch testing ongoing.
    • Calibration (Perley): Gain compression/expansion ongoing (Hayward).
    • Brentjens: Holography readiness/status.
  • WIDAR (Rupen):
    • Notes (Rupen/Sowinski): BlB's fully populated; StB rack fully populated (less deformatters). Data rate work now supports the most challenging RSRO programs (just over 25 MB/s). Pushing to faster dump times (120 ms sustained for OSRO configurations). Pushing to improve configuration times, resolve remaining missing BDFs (1-2 times/night).
  • Computing (Robnett):
    • Notes: Issues with lustre filling.

Software status and updates (Butler)

Commissioning Group Reports and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!

Commissioning priorities for the week (McMullin, Chandler)

Configuration move week (work around these activities:
  • Mon
    • 1000-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Tue: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations and testing
  • Wed:
    • 0930-1730: 11A-229
    • 1800-0900: Science operations and testing
  • Thu:
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Fri (not-yet-a-holiday):
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Sat-Sun
    • Science operations + planned testing

Issues Affecting Science Observing:

  • Quick summary:
    • Good weekend; stabilized data acquisition software issues for the weekend. Some outstanding development items below; some suspected hardware issues, some puzzles.

  • Correlator: [Rupen/Carlson w/ Momjian/Sault] V stokes issues.Identified possible issues in the noise related to the 3-level phase rotator. Water maser data taken.
  • Correlator: [Pokorny/Rupen/McMullin] time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags). .Needs close-out testing.
  • Correlator: [Clark/Carlson] Delay rocking across bands. Ongoing discussion of implementation; pending Bryan's return
    • Array: [All] Zoo of other effects (some data are very good, some still seeing antenna-based IF/Polarization issue).
    • Array: Hunter: OPT testing revealed some oddities with North Arm antennas (status)?
    • Array: Chandler: Stepped (asymptotic) gain changes on some antenna IFs (status)?
    • Array: Hughes: Rapid phase changes at L band
    • Array: EA17 IFC amplitude jumps
    • Array: EA15 B/D not fringing
    • Array: EA10 SPW9 IF B/D phase drifts

Testing This week/soon

  • [Sowinski et al]; Executor update testing
    • Phased array testing
  • [Pokorny et al]; CBE data rate testing (fast dumps); configuration timings
  • Planetary tracking tests
  • Holography scanning tests
  • [Goss et al]; Sub-band edge effects testing.
  • [Owen]: L band deep field study
  • [Carlson et al]; V Stokes follow-up.
  • Subarrays: cleanup.
  • 3-bit sampler testing; automatic gain setting.
  • Stage 2 mixer testing to relax sub 128 MHz BW grids
  • astrometry mode testing.
  • Correlator: Rupen: S-band shadow testing
  • Spectral line: Roy/Dhawan: HI High resolution bandpass testing
  • Baseband edge/filter rolloff.
  • Sub-band dependent fshifts.
  • Array: *; Individual issue sleuthing.

Commissioning Groups

    • Adv: Low band (Owen):
      • Notes:
    • Adv: Astrometry (Mioduszewski):
      • Notes: Preliminary results from Vivek/Andreas. -- follow-up
  • Calibrator Models
    • Notes: Reduction underway.
  • HPC (Bhatnagar/Owen/Golap):
    • Notes: ~500 GB data set obtained; status; talk this Thursday at 1000.
  • Imaging (Ott/Golap):
    • Notes:
  • Algorithm (Bhatnagar, Owen):

RSRO/ECSO Reports

RSRO A configuration status: https://safe.nrao.edu/wiki/bin/view/EVLA/2011BRSROStatus

Individual reports (updates)

Misc:

Demo Science Status

AOB

-- JosephMcMullin - 2011-08-16
Topic revision: r1 - 2011-08-16, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback